OpenStudy (anonymous):

2x + y = -2 x + y = 5 The x-coordinate of the solution to the system shown is _____.

OpenStudy (anonymous):

Sounds like you need to solve for x. I would use subtraction for this system of equations.

OpenStudy (texaschic101):

I would use substitution

OpenStudy (texaschic101):

lol...MagicallyFarting.....what a name

OpenStudy (anonymous):

2x+y=-2 - x+y=5 --------- x=-7

OpenStudy (anonymous):

Lol :D Don't see a name like that everyday. Haha.

OpenStudy (texaschic101):

so true

OpenStudy (anonymous):

No you do not haha

OpenStudy (texaschic101):

x + y = 5 y = 5 - x now sub 5 - x in for y in the 1st equation

OpenStudy (texaschic101):

can you do that Zachary ?

OpenStudy (anonymous):

x+5-xy=5?

OpenStudy (texaschic101):

2x + y = -2 (sub 5 - x in for y) 2x + 5 - x = -2 x + 5 = -2 x = -2-5 x = -7 understand ?

OpenStudy (anonymous):

ohhhhh yea now i do

OpenStudy (texaschic101):

if you wanted to solve for y, you would sub -7 in for x in the 2nd equation and solve for y